Jerry deposited $160 at the bank in one-dollar and five-dollars bills.He deposited three times as many one-dollar bills as five-dollars bills did Jerry deposit at the bank?
Found 2 solutions by priya, sarah_adam:
Answer by priya(51)

In word problems first try to form equations with the informations given in the question.
Let X be the number of one dollar bills and Y be the number of five dollar bills.
then,
EQ 1. 1*X + 5*Y = 160( TOTAL AMOUNT)

EQ 2 . 1 * X = 3 * 5Y
X = 15Y
substitute X = 15Y IN EQ 1.,
we get 15Y + 5Y = 160,
20Y = 160
therefore Y= 160/20 = 8
put Y = 8 in eq 2.,we get
X = 15 * 8
= 120.
the answer is, the number of one -dollar bill is 120 and the number of five -dollar bills is 8..


WANT TO LEARN ONLINE..VISIT MY SITE!!!
PRIYA

Answer by sarah_adam(201)

Let the number of one dollar bills be x
so x+5y = 160
but given x = 3*(5y)= 15y
15y+5y = 160
20y = 160
y = 8
x = 15(8) = 120
number of 5 dollar bills : 8
number of 1 dollar bills : 120